# Imitation from Diverse Behaviors: Wasserstein Quality Diversity Imitation Learning with Single-Step Archive Exploration

### v1

@@ 4.3 Performance
Table 1: Ablation study results on the effect of latent Wasserstein adversarial training (-WGAIL).
HalfCheetah Walker2d Humanoid
mCWAE-WGAIL-Bonus $\mathbf{3.80\times 10^{6}}$ 98.28 4,553 1,547 $\mathbf{3.07\times 10^{6}}$ 87.39 4,142 1,407 $\mathbf{6.40\times 10^{6}}$ 75.05 6,875 3,447
mCWAE-GAIL-Bonus $3.07\times 10^{7}$ 99.32 3,350 1,235 $1.89\times 10^{6}$ 56.88 1,860 1,337 $3.26\times 10^{6}$ 55.95 5,182 2,530
WAE-WGAIL $3.74\times 10^{6}$ 93.77 5,463 1,615 $2.11\times 10^{6}$ 73.20 3,327 1,148 $3.98\times 10^{6}$ 68.08 8,553 2,359
WAE-GAIL $2.91\times 10^{6}$ 92.85 3,792 1,178 $1.21\times 10^{6}$ 43.59 1,357 1,097 $4.01\times 10^{6}$ 66.45 5,721 2,381
Table 2: Ablation study results on the effects of single-step archive exploration (-Bonus) and measure conditioning (mC).
HalfCheetah Walker2d Humanoid
mCWAE-WGAIL-Bonus $\mathbf{3.80\times 10^{6}}$ 98.28 4,553 1,547 $\mathbf{3.07\times 10^{6}}$ 87.39 4,142 1,407 $\mathbf{6.40\times 10^{6}}$ 75.05 6,875 3,447
WAE-WGAIL-Bonus $2.66\times 10^{7}$ 76.89 4,788 1,246 $2.55\times 10^{6}$ 73.93 3,310 1,381 $4.97\times 10^{6}$ 67.27 8,620 3,09
mCWAE-WGAL $3.78\times 10^{7}$ 87.76 5,728 1,623 $2.37\times 10^{6}$ 71.61 3,611 1,326 $3.54\times 10^{6}$ 66.96 8,485 2,114
WAE-WGAIL $3.74\times 10^{6}$ 93.77 5,463 1,615 $2.11\times 10^{6}$ 73.20 3,327 1,148 $3.98\times 10^{6}$ 68.08 8,553 2,359
@@ 4.4 Ablations
4.4.1 Ablation on Latent Wasserstein Adversarial Training
To study the effect of latent Wasserstein adversarial training, we additionally introduce mCWAE-GAIL-Bonus, which reinforces WAE-GAIL with measure conditioning and a single-step archive exploration bonus. Table 1 compares the QD performance of our mCWAE-WGAIL-Bonus with mCWAE-GAIL-Bonus and WAE-WGAIL. As we can see, applying Wasserstein adversarial training improves the QD-Score of WAE-GAIL and mCWAE-GAIL-Bonus by 27.5% and 23.7% respectively on HalfCheetah. In Walker2d, the latent Wasserserstein adversarial training enhances the QD-Score of WAE-GAIL and mCWAE-GAIL-Bonus by 74.3% and 62.4%, which is significant. Moreover in Humanoid, mCWAE-WGAIL-Bonus achieves 2x QD-Score of mCWAE-GAIL-Bonus, impressively outperforming the expert (i.e., PPGA-trueReward) by 12%. The learning curves are shown in Figure 6 in the Appendix D.2.
4.4.2 Ablations on Single-Step Archive Exploration and Measure Conditioning
Table 2 compares the effects of single-step archive exploration bonus and measure conditioning. Measure conditioning improves the QD-Score for WAE-WGAIL-Bonus by 42.8%. In Walker2d, the contributions of measure conditioning to WAE-WGAL and WAE-WGAIL-Bonus are 12.3% and 20.3% respectively. In Humanoid, mCWAE-WGAIL-Bonus outperforms WAE-WGAIL-Bonus by 28.7%. Now, we study the effect of the single-step archive exploration bonus. The bonus does not show much improvement on HalfCheetah. Comparing mCWAE-WGAIL-Bonus and mCWAE-WGAIL, we found that the contributions of bonus on the QD-Score improvement are 29.5% in Walker2d and 80% in Humanoid. Comparing WAE-WGAIL-Bonus and WAE-WGAIL, the contributions of the bonus on the QD-Score improvement are 20.8% in Walker2d and 24.8% in Humanoid. We illustrate the corresponding learning curves of this ablation study in Figure 7 of Appendix D.2.
5 Conclusion
@@ Appendix A Algorithms
Algorithm 5 WAE-GAIL.
@@
Algorithm 6 WAE-WGAIL.
@@ Appendix B Hyperparameter Setting
B.1 Hyperparameters for PPGA
Table 3 summarizes a list of hyperparameters for PPGA policy updates.
Table 3: List of relevant hyperparameters for PPGA shared across all environments.
@@ B.1 Hyperparameters for PPGA
B.2 Hyperparameters for IL
Table 4 summarizes a list of hyperparameters for AIRL, GAIL, measure-conditioned GAIL, and MConbo-GAIL.
Table 4: List of relevant hyperparameters for GAIL and WAEs shared across all environments.
@@ B.2 Hyperparameters for IL
Table 5 summarizes a list of hyperparameters for GIRIL.
Table 5: List of relevant hyperparameters for GIRIL shared across all environments.
@@ Appendix C Demonstration Details
Table 6: Demonstrations statistics on the three environments.
@@ Appendix D Additional Experimental Results
D.1 Quantitative Overall Performance.
Table 7: Compare the QD performance of our mCWAE-WGAIL-Bonus against WAE-WGAIL, GAIL, PWIL, AIRL, MaxEntIRL and GIRIL across three environments. Cov, Best, and Avg refer to Coverage, Best Reward, and Average Reward respectively.
@@
mCWAE-WGAIL-Bonus $3.80\times 10^{6}$ 98.28 4,553 1,547 $\mathbf{3.07\times 10^{6}}$ 87.39 4,142 1,407 $\mathbf{6.40\times 10^{6}}$ 75.05 6,875 3,447
WAE-WGAL $3.74\times 10^{6}$ 93.77 5,463 1,615 $2.11\times 10^{6}$ 73.20 3,327 1,148 $3.98\times 10^{6}$ 68.08 8,553 2,359
GAIL $2.02\times 10^{6}$ 67.83 5,115 1,165 $2.48\times 10^{6}$ 69.29 4,031 1,429 $1.86\times 10^{6}$ 82.27 6,278 924
PWIL $\mathbf{4.17\times 10^{6}}$ 99.65 3,626 1,675 $2.27\times 10^{6}$ 64.45 2,835 1,410 $1.12\times 10^{6}$ 91.73 841 492
AIRL $3.11\times 10^{6}$ 83.57 5,183 1,410 $2.53\times 10^{6}$ 70.53 4,280 1,437 $2.31\times 10^{6}$ 71.47 7,661 1,308
MaxEntIRL $1.12\times 10^{6}$ 85.48 2,570 523 $1.80\times 10^{6}$ 68.81 3,756 1,046 $1.82\times 10^{6}$ 83.27 4,658 882
GIRIL $2.17\times 10^{6}$ 95.96 3,466 909 $0.52\times 10^{6}$ 25.08 1,139 821 $4.33\times 10^{6}$ 67.40 6,992 2,590
D.2 Learning Curves of Ablation Studies
@@
Figure 6: Learning curve comparison of our mCWAE-WGAIL-Bonus against WAE-GAIL-Bonus, WAE-WGAIL and WAE-GAIL.
@@
Figure 7: Learning curve comparison of our mCWAE-WGAIL-Bonus against the components: WAE-WGAIL-Bonus, mCWAE-WGAIL and WAE-WGAIL.

### v2

@@ 4.4. Ablations
4.4.1. Ablation on Latent Wasserstein Adversarial Training
To study the effect of latent Wasserstein adversarial training, we additionally introduce mCWAE-GAIL-Bonus, which reinforces WAE-GAIL with measure conditioning and a single-step archive exploration bonus. Table 7 in Appendix D.2 compares the QD performance of our mCWAE-WGAIL-Bonus with mCWAE-GAIL-Bonus and WAE-WGAIL. As we can see, applying Wasserstein adversarial training improves the QD-Score of WAE-GAIL and mCWAE-GAIL-Bonus by 27.5% and 23.7% respectively on HalfCheetah. In Walker2d, the latent Wasserserstein adversarial training enhances the QD-Score of WAE-GAIL and mCWAE-GAIL-Bonus by 74.3% and 62.4%, which is significant. Moreover in Humanoid, mCWAE-WGAIL-Bonus achieves 2x QD-Score of mCWAE-GAIL-Bonus, impressively outperforming the expert (i.e., PPGA-trueReward) by 12%. The learning curves are shown in Figure 6 in Appendix D.2.
4.4.2. Ablations on Single-Step Archive Exploration and Measure Conditioning
Table 8 in of Appendix D.2 compares the effects of single-step archive exploration bonus and measure conditioning. Measure conditioning improves the QD-Score for WAE-WGAIL-Bonus by 42.8%. In Walker2d, the contributions of measure conditioning to WAE-WGAL and WAE-WGAIL-Bonus are 12.3% and 20.3% respectively. In Humanoid, mCWAE-WGAIL-Bonus outperforms WAE-WGAIL-Bonus by 28.7%. Now, we study the effect of the single-step archive exploration bonus. The bonus does not show much improvement on HalfCheetah. Comparing mCWAE-WGAIL-Bonus and mCWAE-WGAIL, we found that the contributions of bonus on the QD-Score improvement are 29.5% in Walker2d and 80% in Humanoid. Comparing WAE-WGAIL-Bonus and WAE-WGAIL, the contributions of the bonus on the QD-Score improvement are 20.8% in Walker2d and 24.8% in Humanoid. We illustrate the corresponding learning curves of this ablation study in Figure 7 of Appendix D.2.
5. Conclusion
@@
{acks}
This project is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority. Zhenglin Wan contributed to this publication whilst serving his internship with A*STAR CFAR under the Foreign Student Attachment scheme.
@@ Appendix A Algorithms
Algorithm 4 WAE-GAIL reward model updates.
@@
Algorithm 5 WAE-WGAIL reward model updates.
@@
Algorithm 6 mCWAE-WGAIL reward model updates.
2:Initialize the parameters of the encoder $Q_{\phi}$,
4:Method: Update reward model in Algorithm 1 (step 17)
5:def update($\mathcal{D},\pi_{\theta}$):
10: Update $D_{\mathcal{W}}$ by ascending:
11: Update $Q_{\phi}$ and $G_{\psi}$ by descending:
12: Repeat until the model converges or the number of epochs is reached
13:return Updated $Q_{\phi}$, $G_{\psi}$ and $D_{\mathcal{W}}$
@@ Appendix B Hyperparameter Setting
B.1. Hyperparameters for PPGA
Table 1 summarizes the list of hyperparameters for PPGA policy updates.
Table 1. List of relevant hyperparameters for PPGA shared across all environments.
@@ B.1. Hyperparameters for PPGA
B.2. Hyperparameters for IL
Table 2 summarizes the list of hyperparameters for AIRL, GAIL, measure-conditioned GAIL, and MConbo-GAIL.
Table 2. List of relevant hyperparameters for GAIL and WAEs shared across all environments.
@@ B.2. Hyperparameters for IL
Table 3 summarizes the list of hyperparameters for GIRIL.
Table 3. List of relevant hyperparameters for GIRIL shared across all environments.
@@ Appendix C Demonstration Details
Table 4. Demonstration statistics on the three environments.
@@ Appendix D Additional Experimental Results
D.1. Overall results: quantitative performance and Tukey HSD test
Table 5. QD performance comparison of our mCWAE-WGAIL-Bonus against WAE-WGAIL, GAIL, PWIL, AIRL, MaxEntIRL and GIRIL across three environments. Cov, Best, and Avg refer to Coverage, Best Reward, and Average Reward respectively. Bold indicates the highest two metric scores.
@@ D.1. Overall results: quantitative performance and Tukey HSD test
mCWAE-WGAIL-Bonus $\mathbf{3.80\times 10^{6}}\pm\mathbf{2.85\times 10^{5}}$ 98.28$\pm$ 0.42 4,553$\pm$547 1,547$\pm$118 $\mathbf{3.07\times 10^{6}}\pm\mathbf{9.75\times 10^{4}}$ 87.39$\pm$4.32 4,142$\pm$496 1,407$\pm$26 $\mathbf{6.40\times 10^{6}}\pm\mathbf{6.31\times 10^{5}}$ 75.05$\pm$4.82 6,875$\pm$254 3,447$\pm$562
WAE-WGAIL $3.74\times 10^{6}\pm 7.55\times 10^{5}$ 93.77 $\pm$4.67 5,463$\pm$136 1,620$\pm$395 $2.11\times 10^{6}\pm 3.25\times 10^{5}$ 73.20$\pm$6.24 3,327$\pm$192 1,148$\pm$99 $3.98\times 10^{6}\pm 3.38\times 10^{5}$ 68.08$\pm$3.9 8,553$\pm$521 2,359$\pm$321
PWIL $\mathbf{4.17\times 10^{6}}\pm\mathbf{4.07\times 10^{5}}$ 99.65$\pm$0.28 3,626$\pm$312 1,682$\pm$165 $2.27\times 10^{6}\pm 1.49\times 10^{5}$ 64.45$\pm$0.66 2,835$\pm$470 1,410$\pm$106 $1.12\times 10^{6}\pm 7.20\times 10^{4}$ 91.73$\pm$4.11 841$\pm$96 492$\pm$24
AIRL $3.11\times 10^{6}\pm 1.71\times 10^{6}$ 83.57$\pm$9.65 5,183$\pm$1,735 1,410$\pm$700 $\mathbf{2.53\times 10^{6}\pm 2.59\times 10^{5}}$ 70.53$\pm$2.3 4,280$\pm$326 1,437$\pm$141 $2.31\times 10^{6}\pm 1.15\times 10^{5}$ 71.47$\pm$7.59 7,661$\pm$705 1,308$\pm$144
MaxEntIRL $1.12\times 10^{6}\pm 3.54\times 10^{5}$ 85.48$\pm$10.65 2594$\pm$680 525$\pm$152 $1.80\times 10^{6}\pm 5.13\times 10^{5}$ 68.83$\pm$0.73 3,756$\pm$318 1,046$\pm$29 $1.82\times 10^{6}\pm 3.12\times 10^{5}$ 83.27$\pm$5.77 4,658$\pm$1489 882$\pm$190
GIRIL $2.17\times 10^{6}\pm 8.85\times 10^{5}$ 95.96$\pm$0.79 3,466$\pm$1018 909$\pm$377 $0.52\times 10^{6}\pm 1.33\times 10^{5}$ 25.08$\pm$3.93 1,139$\pm$25 821$\pm$112 $\mathbf{4.33\times 10^{6}\pm 2.58\times 10^{5}}$ 67.40$\pm$6.05 6,992$\pm$915 2,590$\pm$254
Table 6. Tukey HSD test of the QD Score for the comparisons to baselines. $+$ indicates a positive effect of mCWAE-WGAIL-Bonus. The significant pairwise comparisons ($p$¡=0.05) are boldfaced.
group1 group2 HalfCheetah Walker2d Humanoid
AIRL mCWAE-WGAIL-Bonus + 0.9825 + 0.2094 + 0.0
GAIL mCWAE-WGAIL-Bonus + 0.4375 + 0.1432 + 0.0
GIRIL mCWAE-WGAIL-Bonus + 0.5333 + 0.0 + 0.0007
MaxEntIRL mCWAE-WGAIL-Bonus + 0.0928 + 0.0005 + 0.0
PWIL mCWAE-WGAIL-Bonus - 0.9992 + 0.0241 + 0.0
D.2. Ablation Study: quantitative performance, learning curves, and Tukey HSD test
Table 7. Ablation study results on the effect of latent Wasserstein adversarial training (-WGAIL). Bold indicates the highest metric score.
HalfCheetah Walker2d Humanoid
mCWAE-WGAIL-Bonus $\mathbf{3.80\times 10^{6}}\pm\mathbf{2.85\times 10^{5}}$ 98.28$\pm$ 0.42 4,553$\pm$547 1,547$\pm$118 $\mathbf{3.07\times 10^{6}}\pm\mathbf{9.75\times 10^{4}}$ 87.39$\pm$4.32 4,142$\pm$496 1,407$\pm$26 $\mathbf{6.40\times 10^{6}}\pm\mathbf{6.31\times 10^{5}}$ 75.05$\pm$4.82 6,875$\pm$254 3,447$\pm$562
mCWAE-GAIL-Bonus $3.07\times 10^{7}\pm 4.03\times 10^{5}$ 99.32$\pm$0.64 3,350$\pm$570 1,235$\pm$158 $1.89\times 10^{6}\pm 2.11\times 10^{5}$ 56.88$\pm$2.39 1,860$\pm$310 1,337$\pm$178 $3.26\times 10^{6}\pm 3.03\times 10^{5}$ 55.95$\pm$17.41 5,182$\pm$395 2,542$\pm$639
WAE-WGAIL $3.74\times 10^{6}\pm 7.55\times 10^{5}$ 93.77 $\pm$4.67 5,463$\pm$136 1,620$\pm$395 $2.11\times 10^{6}\pm 3.25\times 10^{5}$ 73.20$\pm$6.24 3,327$\pm$192 1,148$\pm$99 $3.98\times 10^{6}\pm 3.38\times 10^{5}$ 68.08$\pm$3.9 8,553$\pm$521 2,359$\pm$321
WAE-GAIL $2.91\times 10^{6}\pm 2.02\times 10^{6}$ 92.85$\pm$10.05 3,792$\pm$1,227 1,178$\pm$794 $1.21\times 10^{6}\pm 3.98\times 10^{5}$ 43.59$\pm$13.15 1,357$\pm$147 1,097$\pm$58 $4.01\times 10^{6}\pm 1.15\times 10^{6}$ 66.45$\pm$10.99 5,721$\pm$847 2,381$\pm$347
Table 8. Ablation study results on the effects of single-step archive exploration (-Bonus) and measure conditioning (mC). Bold indicates the highest metric score.
HalfCheetah Walker2d Humanoid
mCWAE-WGAIL-Bonus $\mathbf{3.80\times 10^{6}}\pm\mathbf{2.85\times 10^{5}}$ 98.28$\pm$ 0.42 4,553$\pm$547 1,547$\pm$118 $\mathbf{3.07\times 10^{6}}\pm\mathbf{9.75\times 10^{4}}$ 87.39$\pm$4.32 4,142$\pm$496 1,407$\pm$26 $\mathbf{6.40\times 10^{6}}\pm\mathbf{6.31\times 10^{5}}$ 75.05$\pm$4.82 6,875$\pm$254 3,447$\pm$562
WAE-WGAIL-Bonus $2.66\times 10^{7}\pm 2.01\times 10^{6}$ 76.89$\pm$15.46 4,788$\pm$1,339 1,246$\pm$700 $2.55\times 10^{6}\pm 1.28\times 10^{5}$ 73.93$\pm$2.58 3,310$\pm$114 1,381$\pm$99 $4.97\times 10^{6}\pm 7.55\times 10^{5}$ 67.27$\pm$12.64 8,620$\pm$732 3,099$\pm$826
mCWAE-WGAIL $3.79\times 10^{7}\pm 1.89\times 10^{6}$ 87.76$\pm$12.95 5,728$\pm$1429 1,623$\pm$693 $2.37\times 10^{6}\pm 9.3\times 10^{5}$ 71.61$\pm$2.0 3,611$\pm$524 1,326$\pm$73 $3.54\times 10^{6}\pm 3.09\times 10^{5}$ 66.96$\pm$0.69 8,485$\pm$290 2,114$\pm$194
WAE-WGAIL $3.74\times 10^{6}\pm 7.55\times 10^{5}$ 93.77 $\pm$4.67 5,463$\pm$136 1,620$\pm$395 $2.11\times 10^{6}\pm 3.2\times 10^{5}$ 73.20$\pm$6.24 3,327$\pm$192 1,148$\pm$99 $3.98\times 10^{6}\pm 3.38\times 10^{5}$ 68.08$\pm$3.9 8,553$\pm$521 2,359$\pm$321
@@ D.2. Ablation Study: quantitative performance, learning curves, and Tukey HSD test
Figure 6. Learning curve comparison of our mCWAE-WGAIL-Bonus against WAE-GAIL-Bonus, WAE-WGAIL and WAE-GAIL.
Table 9. Tukey HSD test of the QD Score for the ablation study on Wasserstein adversarial training. $+$ indicates a positive effect of mCWAE-WGAIL-Bonus. The significant pairwise comparisons ($p$¡=0.05) are boldfaced.
group1 group2 HalfCheetah Walker2d Humanoid
GAIL mCWAE-WGAIL-Bonus + 0.4844 + 0.3009 + 0.0003
WAE-GAIL mCWAE-WGAIL-Bonus + 0.9127 + 0.0005 + 0.027
WAE-WGAIL mCWAE-WGAIL-Bonus + 1.0 + 0.0433 + 0.0256
mCWAE-GAIL-Bonus mCWAE-WGAIL-Bonus + 0.9533 + 0.0133 + 0.0051
@@
Figure 7. Learning curve comparison of our mCWAE-WGAIL-Bonus against WAE-WGAIL-Bonus, mCWAE-WGAIL and WAE-WGAIL.
Table 10. Tukey HSD test of the QD Score for the ablation study on bonus and measure conditioning. $+$ indicates a positive effect of mCWAE-WGAIL-Bonus. The significant pairwise comparisons ($p$¡=0.05) are boldfaced.
group1 group2 HalfCheetah Walker2d Humanoid
WAE-WGAIL mCWAE-WGAIL-Bonus + 1.0 + 0.004 + 0.0088
WAE-WGAIL-Bonus mCWAE-WGAIL-Bonus + 0.8562 + 0.0886 + 0.1079
mCWAE-WGAIL mCWAE-WGAIL-Bonus + 1.0 + 0.0239 + 0.0032